On a farmer’s market in Essex, the demand and supply functions for (a kilogram (kg) of) oranges (1), (a kg of) lemon (2) and (a kg of) nectarines (3) are:
q_1^d=-2p_1-4p_2+2p_3+160,q_1^s=2p_1+80
q_2^d=-4p_1-2p_2+5p_3+200,q_2^s=3p_2+160
q_3^d=2p_1+4p_2-5p_3+300,q_3^s=p_3+90
Use the equilibrium condition q_i^d=q_i^s to rewrite this system in the form Ap=d, where A is a 3×3 matrix of coefficients, p=(■(p_1@p_2@p_3 )), and d is a 3×1 vector. Then, use Cramer’s rule to solve for p_1 and p_3 (ONLY solve for p_1 and p_3!)
1
Expert's answer
2020-06-23T19:51:58-0400
"-2p_1-4p_2+2p_3+160=2p_1+80""-4p_1-2p_2+5p_3+200=3p_2+160""2p_1+4p_2-5p_3+300=p_3+90"


"-4p_1-4p_2+2p_3=-80""-4p_1-5p_2+5p_3=-40""2p_1+4p_2-6p_3=-210"

"A=\begin{pmatrix}
 -4 & -4 & 2 \\
 -4 & -5 & 5 \\
 2 & 4 & -6
\end{pmatrix}"

"p=\begin{pmatrix}
 p_1 \\
 p_2 \\
p_3
\end{pmatrix}, d=\begin{pmatrix}
 -80 \\
 -40 \\
-210
\end{pmatrix}"

"Ap=d"

"\Delta=\begin{vmatrix}
 -4 & -4 & 2 \\
 -4 & -5 & 5 \\
 2 & 4 & -6
\end{vmatrix}="

"=-4\begin{vmatrix}
 -5 & 5 \\
 4 & -6
\end{vmatrix}-(-4)\begin{vmatrix}
 -4 & 5 \\
 2 & -6
\end{vmatrix}+2\begin{vmatrix}
 -4 & -5 \\
 2 & 4
\end{vmatrix}="

"=-4(30-20)+4(24-10)+2(-16+10)=4\not=0"

Solutions exist.


"\Delta_1=\begin{vmatrix}
 -80 & -4 & 2 \\
 -40 & -5 & 5 \\
 -210 & 4 & -6
\end{vmatrix}="

"=-80\begin{vmatrix}
 -5 & 5 \\
 4 & -6
\end{vmatrix}-(-4)\begin{vmatrix}
 -40 & 5 \\
 -210 & -6
\end{vmatrix}+2\begin{vmatrix}
 -40 & -5 \\
 -210 & 4
\end{vmatrix}="

"=-80(30-20)+4(240+1050)+2(-160-1050)=1940"


"\Delta_3=\begin{vmatrix}
 -4 & -4 & -80 \\
 -4 & -5 & -40 \\
 2 & 4 & -210
\end{vmatrix}="

"=-4\begin{vmatrix}
 -5 & -40 \\
 4 & -210
\end{vmatrix}-(-4)\begin{vmatrix}
 -4 & -40 \\
 2 & -210
\end{vmatrix}-80\begin{vmatrix}
 -4 & -5 \\
 2 & 4
\end{vmatrix}="

"=-4(1050+160)+4(840+80)-80(-16+10)=-680"


"p_1={1940\over 4}=485"

"p_3={-680\over 4}=-170"